SYNCHROTRON FAR INFRARED SPECTROSCOPY: HIGHER RESOLUTION AND LONGER WAVE-LENGTHS AT THE CANADIAN LIGHT SOURCE

\underline{A.R.W. McKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, Canada; D.R.T. APPADOO^{a}, Canadian Light Source, 101 Perimeter Road, University of Saskatchewan, Saskatoon, SK S7N 0X4, Canada.

In order to take full advantage of the high spectral resolution capability of the Bruker IFS125 Fourier transform spectrometer (MOPD = 9.4 m) at the Canadian Light Source far infrared beamline, we have interfaced a large 2 m gas cell which allows longer absorption paths, lower sample pressures, and hence reduced pressure broadening. We are using acrolein (CH_{2}CHCHO) for testing. Interesting in its own right, this molecule is also of importance in astrophysics, combustion chemistry, and human respiration (cigarette smoke, smog). Our initial CLS work^{b} on ν_{12} (∼564 cm^{−1}) and ν_{17} (∼593 cm^{−1}) has now been extended to ν_{18} (∼158 cm^{−1}), 2ν_{18} (∼314 cm^{−1}), ν_{13} (∼323 cm^{−1}), 3ν_{18} (∼469 cm^{−1}), and further vibrational levels. We experience significant gains in source intensity and signal-to-noise ratio due to the brightness advantage of synchrotron radiation compared to conventional thermal far infrared sources, though source noise (due to mechanical vibrations) remains a concern. Even greater gains at long wavelengths (<50 cm^{−1}) may be possible using coherent synchrotron radiation (CSR), achieved by shortening the electron bunch length in the synchrotron storage ring.^{c} Initial testing of CSR at CLS has already yielded a recognizable high-resolution spectrum of N_{2}O pure rotational transitions in the 10 - 20 cm^{−1}range.

aCurrent address: Australian Synchrotron, 800 Blackburn Rd, Clayton VIC 3168, Australia.

bA.R.W. McKellar, D.W. Tokaryk, L.-H. Xu, D.R.T. Appadoo, and T. May, J. Mol. Spectrosc. 241, 31 (2007).

cM. Abo-Bakr, J. Feikes, K. Holldack, P. Kuske, W. B. Peatman, U. Schade, G. Wüstefeld, and H.-W. Hübers, Phys. Rev.

Letters 90, 094801 (2003).